• Total Output: R 30
• Total Income: R 30
• Consumption: R 22
• Government Spending: R 6
• Total tax revenue: R 5.5
• Transfer payments: R 1.5
Use the above data to calculate:
a. disposable income, savings and net taxes.
b. Is the government running a budget surplus, budget deficit, or balanced budget? Demonstrate.
c. Determine planned investment. (note that investment is equal to the sum of private and public saving)
1
Expert's answer
2020-04-15T09:49:53-0400

O=R30

I=R30

C=R22

G=R6

Tr=R5.5

t=R1.5

a). disposable income,"Y=C+I+G"

Y=R22+R30+R6=R58

Savings, "S=I"

S=R30

Net taxes,"T=I-t-Tr"

T=R30-R1.5-R5.5=R23

b). The government running a Budget deficit.

Tr<G

c).  planned investment,"pI=S=I" =R30